a dirt path 4 feet width surrounds a rectangular garden that is 38 feet longer and 17 feet wide. find the area of the dirt path?
Answer by mananth(16949)

Length of garden 38 feet
Width = 17 feet
width of dirtpath= 4 feet
Area = 38 * 17 646 ft^2
Length of garden +width of dirtpath 38 + 4
Width of garden + width of dirtpath 17 + 4
Area ofgarden + dirt path=(38+4 )(17+4)
Area ofgarden + dirt path=42*21
Area ofgarden + dirt path=882
Area of walkway = area of (pool+walkway)-areaof garden= 882 - 646
Area of dirt Path=236 ft^2